OpenStudy (anonymous):

*Multiple Choice! Just need to make sure I'm right! Will give medal to first reply!* Jess plans to increase the number of hours she spends doing exercises. She can increase the hours in the following ways: Month: 1 2 3 4 5 Option 1: 0.5 0.8 1.1 1.4 1.7 Option 2: 0.5 1 2.0 4.0 8.0 Option 3: 0.5 0.55 0.61 0.67 0.74

OpenStudy (anonymous):

If Jess wants to increase the hours linearly, which option(s) should she choose? Only option 1, because it shows equal increases in equal intervals of time. Only option 2, because it shows equal increases in equal intervals of time. Either option 1 or option 2, because they show increases in time by the same percentage. Either option 2 or option 3, because they show increases in time by the same percentage.

OpenStudy (anonymous):

I think it's the third choice

OpenStudy (anonymous):

The Answer is Option 1, 0.5 + 0.3 = 0.8+0.3 = 1.1 + 0.3 = 1.4 +0.3 =1.7 It adds 0.3 every time.

OpenStudy (anonymous):

Oh, okay. Yea, I see that. Thanks! @LovenEpic